OpenStudy (anonymous):

Solve the system, using substitution. Medal for correct steps and for correct ordered pair. x = 4y and 2x + 3y = 44

hero (hero):

2(4y) + 3y = 44 8y + 3y = 44 11y = 44 y = 44/11 y = 4

OpenStudy (anonymous):

Hope this helps :)

hero (hero):

How does y = 11? you wrote 11y = 44 if y = 11 then 11*11 = 44 121 = 44 That is not true

hero (hero):

y = 4 x = 16

hero (hero):

You clearly do not understand what you are doing wrong.

OpenStudy (anonymous):

you have two variables in both equations, which makes it hard to solve. thus, you want to substitute 4y into the second equation for x, so the second equation only has the variable y. you can do this because the first equation states that x is equal to 4y

OpenStudy (anonymous):

2(4y) + 3y = 44 8y + 3y = 44 11y = 44 y = 4

OpenStudy (anonymous):

now, plug that into the first equation, x = 4y, to find x x = 4(4) x = 16 so your answer is x = 16, y =4, or (16, 4)

OpenStudy (anonymous):

Am i right now Mr. Hero?

hero (hero):

I'm pretty sure the original poster ran away with the wrong answer. It's too late now. What you should do next time is if you see someone post a correct answer, just give them the medal or say "I agree" and move on. When you go behind someone and post a different answer, it confuses the OP.

OpenStudy (anonymous):

No not like that

OpenStudy (anonymous):

2x - 3y = 4 x + 4y = -9 i need major help with this. Anyone?